OpenStudy (anonymous):

what is the prime factorization of 75^75?

OpenStudy (anonymous):

\[75^{75}\]

OpenStudy (anonymous):

the better question is, what is the prime factorization of 75 ( just ignore the exponent)

OpenStudy (anonymous):

5 times 5 times 3

OpenStudy (anonymous):

ok now since it was originally to the 75th power that means all the prime factorization of 75 are all to the 75th power

OpenStudy (anonymous):

so the answer is\[5^{75} \times 5^{^{75}} \times 3^{75}\]?

OpenStudy (anonymous):

pretty much i mean you could factor out 5^75 as 5*5*5*5*5*5 .... 75 times

OpenStudy (anonymous):

lol okay thanks alot man

jimthompson5910 (jim_thompson5910):

You could combine the two \(\large 5^{75}\) terms to get \(\large 5^{75} \times 5^{75} = 5^{75+75} = 5^{150}\)

OpenStudy (anonymous):

wow i didn't see that thanks

jimthompson5910 (jim_thompson5910):

np